# On the future cover of a sofic shift

### v2

@@ 2.5. The terminal component in the future cover of an irreducible sofic shift
The existence of $\phi$ follows from Corollary 2.24, Proposition 2.22 and Theorem 2.11. The uniqueness follows here by the same argument which was used by Krieger in [Kr1]: The elements of $Y$ that contain synchronized words arbitrarily far ’to the left’ have unique pre-image under $L_{G}$ because they are magic for $(G,L_{G})$ by Exercise 3.3.3 in [LM] or Lemma 4.3 in [MMTW]. (These elements were called $F$-finitary in [Kr1].) Since $G$ is strongly connected it follows from Proposition 3.3.16 in [LM] that they are dense in $X_{G}$. Thus the elements $x\in X_{G}$ for which $L_{G}^{-1}(L_{G}(x))=\{x\}$ are dense in $X_{G}$. Denote this set by $\mathcal{X}_{G}$. If $\phi:X_{G}\to X_{H}$ is a conjugacy such that (10) commutes, $\phi(\mathcal{X}_{G})=\mathcal{X}_{H}$. If $\phi_{1}$ and $\phi_{2}$ are both conjugacies that make the diagram (10) commute, and $x\in\mathcal{X}_{G}$, we have that $\phi_{i}(x)\in\mathcal{X}_{H}$ and
$L_{H}(\phi_{1}(x))=\psi(L_{G}(x))=L_{H}(\phi_{2}(x)),$
implying that $\phi_{1}(x)=\phi_{2}(x)$. Since $\mathcal{X}_{G}$ is dense in $X_{G}$, it follows that $\phi_{1}=\phi_{2}$. ∎
@@ 3. Additional properties of the future cover
Let $(H,L_{H})$ be a labeled graph. A subset $U\subseteq V_{H}$ of vertexes in $H$ is a hereditary subset when $e\in E_{H},\ s_{H}(e)\in U\ \Rightarrow\ s_{H}(e)\in U$. When this holds we let $H_{U}$ denote the graph with $V_{H_{U}}:=U$ and $E_{H_{U}}:=\left\{e\in E_{H}:\ t_{H}(e)\in U\right\}$. By restriction the labeling $L_{H}$ defines a labeling of $H_{U}$ and we denote the resulting labeled graph by $(H_{U},L_{H})$. We call this a heritary labeled subgraph of $(H,L_{H})$. Lemma 2.18 can then be improved to the following.
@@
Assume that $(G,L_{G})$ is a right-resolving and regular presentation of the sofic shift $Y$. It follows that there is a labeled graph homomorphism $\theta:(G,L_{G})\to(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ such that
$\theta\left(\mathcal{R}(L_{G})\right)\subseteq\mathcal{R}(L_{\mathbb{K}(Y)}),$
and $\theta(V_{G})$ is a hereditary subset of $V_{\mathbb{K}(Y)}$. If, furthermore, $(G,L_{G})$ is also follower-separated, $\theta$ is a labeled graph isomorphism of $(G,L_{G})$ onto the hereditary labeled subgraph
$(\mathbb{K}(Y)_{\theta(V_{G})},L_{\mathbb{K}(Y)}).$
@@
By Lemma 2.18 and its proof there is a labeled-graph homomorphism $\theta:(G,L_{G})\to(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ such that $\theta(v)=F(L_{G}(x))$ where $x\in X_{G}$, $t_{G}(x_{(-\infty,-1]})=v$ and $f_{G}(v)=F(L_{G}(x))$. Let $F(\theta(v))\overset{a}{\to}F(z)$ be an arrow in $\mathbb{K}(Y)$. Then $a$ is the first letter in an element $w\in F(L_{G}(x)))=f_{G}(v)$, and there is therefore a right-infinite path $\gamma$ in $G$ such that $s_{G}(\gamma)=v$ and $L_{G}(\gamma)=w$. The first arrow in $\gamma$ is an arrow $e\in E_{G}$ with $s_{G}(e)=v$ and $L_{G}(e)=a$. There is therefore an arrow $F(\theta(v))\overset{a}{\to}\theta(t_{G}(e))$ in $\mathbb{K}(Y)$. Since $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ is right-resolving this implies that $F(z)=\theta(t_{G}(e))$, proving that $\theta(V_{G})$ is a hereditary subset of $V_{\mathbb{K}(Y)}$. The remaining assertions follow now from Lemma 2.18.
@@
Example 3.6.
aabbcc1122334433114411
Example 3.7.
aabb0011
3.2. A universal property of the future cover $L_{G}:X_{\mathbb{K}(Y)}\to Y$.
@@
Let $Y$ be an irreducible sofic shift and $(G,L_{G})$ the minimal right-resolving presentation of $Y$. Let $X$ be an irreducible SFT and $\pi:X\to Y$ a right-closing factor map. There is a sliding block code $\rho:X\to X_{G}$ such that
@@
Let $Y$ be an irreducible SFT and $\pi:X\to Y$ a factor map with $X$ an SFT. Then $\pi$ is isomorphic to the minimal right-resolving cover of $Y$ if and only if
@@
The following arguments are all implicitly contained in [BKM]. First, it is well-known from [BKM] and [LM] that the minimal right-resolving cover has all four properties, and hence so has any cover isomorphic to it. So assume that $\pi:X\to Y$ is a cover for which they all hold. We must show that $\pi$ is then isomorphic to the minimal right-resolving cover $L_{G}:X_{G}\to Y$ of $Y$. For this note that there are sliding block codes $\rho:X\to X_{G}$ and $\rho^{\prime}:X_{G}\to X$ such that
@@
both commute. An application of Lemma 9.1.13 in [LM] shows that $\rho$ and $\rho^{\prime}$ both maps doubly transitive points to doubly transitive points, so condition c) implies that $\rho\circ\rho^{\prime}=\operatorname{id}_{X_{G}}$ and $\rho^{\prime}\circ\rho=\operatorname{id}_{X}$, first on doubly transitive points and then all over by continuity. ∎
@@
Let $(H,L_{H})$ be a regular right-resolving presentation of the sofic shift $Y$. For $x\in X_{H}$ and $N\in\mathbb{N}$ there is a $y\in\mathcal{R}(L_{H})$ such that $y_{i}=x_{i}$ for $i\geq-N$.
@@
Let $(H,L_{H})$ be a right-resolving presentation of the sofic shift $Y$. Then $(H,L_{H})$ is regular if and only if $\mathcal{R}(L_{H})$ is dense in $X_{H}$.
@@
In the language of [BKM] which we have adopted above, Proposition 3.13 shows that the future cover is minimal with respect to regular and right-closing factor maps. The regularity condition on $\pi$ can not be removed: The right-resolving labeled graph $(G,L_{G})$ in Figure 7 presents a sofic shift $Y$ whose future cover $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ is obtained by removing the vertex 10 in Figure 7, and $L_{G}$ does not factor through $L_{\mathbb{K}(Y)}$. The future cover is therefore not minimal with respect to arbitrary right-resolving factor maps in general.
@@
Let $x\in\mathcal{R}(\pi)$ and consider an element $y\in\mathbb{U}(\pi^{\prime}(\rho(x))=\mathbb{U}(\pi(x))$. There is then an element $z\in\mathbb{U}(x)$ such that $\pi(z)=y$. Since $\rho(z)\in\mathbb{U}(\rho(x))$ and $\pi^{\prime}(\rho(z))=y$, it follows that $\rho(x)\in\mathcal{R}(\pi^{\prime})$. ∎
@@ 4.1. Subset constructions
for $K\in X^{\#}$. In particular, we get from $(X,\phi)$ a new dynamical system $(X^{\#},\phi^{\#})$. More generally when $(X,\phi)$ and $(Y,\phi^{\prime})$ are dynamical systems and $\pi:X\to Y$ is a map of dynamical systems in the sense that $\pi\circ\phi=\phi^{\prime}\circ\pi$, the set
@@
Consider two conjugacies $\psi^{\prime}:(X,\phi)\to(X^{\prime},\phi^{\prime})$ and $\psi:(Y,\mu)\to(Z,\mu^{\prime})$ of dynamical systems, and factor maps $\pi:(X,\phi)\to(Y,\mu)$ and $\pi^{\prime}:(X^{\prime},\phi^{\prime})\to(Z,\mu^{\prime})$ such that
@@
${X^{\pi}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi^{\prime\#}}$$\scriptstyle{\pi^{\#}}$${X^{\prime\pi^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{\prime\#}}$${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$${Z}$
@@ 4.2. Subset constructions for sofic shifts
To construct an inverse map $\kappa^{-1}:X_{H^{\prime\prime}}\to X_{H}^{L_{H}}$ to $\kappa$, let $x=(x_{i})_{i\in\mathbb{Z}}\in X_{H^{\prime\prime}}$. Since $(H,L_{H})$ is right-resolving, $s_{H^{\prime\prime}}(x_{i})\geq\#s_{H^{\prime\prime}}(x_{i+1})$, and we choose $N\in\mathbb{N}$ such that $s_{H^{\prime\prime}}(x_{i})=\#s_{H^{\prime\prime}}(x_{N})$ for all $i\leq N$. For each $v\in s_{H^{\prime\prime}}(x_{N})$ there is a unique path $\gamma^{v}\in X_{H}$ such that $s_{H}(\gamma^{v}_{N})=v$, $s_{H}(\gamma^{v}_{i})\in s_{H^{\prime\prime}}(x_{i})$ for all $i\in\mathbb{Z}$ and $L_{H}(\gamma^{v})=L_{H^{\prime\prime}}(x)$. Setting
@@
Thus $(X_{G^{\prime\prime}},L_{G^{\prime\prime}})$ is a weakly canonical cover when $(G,L_{G})$ is. However, the conjugacy $\phi^{\prime\prime}$ may not be unique and there is no analogue to the right-inverse for the labeling map as in Krieger’s theorem. We continue therefore to identify a labeled subgraph of $(G^{\prime\prime},L_{G^{\prime\prime}})$ which does give rise to a strongly canonical cover.
@@
$L_{H^{\prime\prime}}\circ\beta_{H}=\operatorname{id}_{Y}.$
@@
Let $y\in Y$ and let $C$ be the component in $H^{\prime\prime}$ for which there is a $N\in\mathbb{Z}$ such that $\beta_{H}(y)_{j}\in C$ for all $j\leq N$. It follows that there is an $M<N$ such that if $z\in X_{H^{\prime\prime}}$ and $L_{H^{\prime\prime}}(z)_{[M,\infty)}=y_{[M,\infty)}$, then $s_{H^{\prime\prime}}(z_{N})\subseteq s_{H^{\prime\prime}}(\beta_{H}(y)_{N})$. (If not, a compactness argument would produce an element $z^{\prime}\in X_{H^{\prime\prime}}$ such that $L_{H^{\prime\prime}}(z^{\prime})=y$ and $s_{H^{\prime\prime}}(z^{\prime}_{N})>\#s_{H^{\prime\prime}}(\beta_{H}(y)_{N})$. A contradiction.) Let $C^{\prime}$ be a component which connects to $C$ and has the largest multiplicity among the components that connect to $C$. Let $\gamma$ be a finite path in $H^{\prime\prime}_{C}$. We can then choose an element $z^{\prime}\in X_{H^{\prime\prime}}$ with the following properties:
@@
    there is a $M^{\prime}<M$ such that $z^{\prime}_{[M^{\prime},M]}$ is a path in $H^{\prime\prime}_{C}$ with $\gamma\subseteq z^{\prime}_{[M^{\prime},M]}$, and
@@
    $z^{\prime}_{[M,\infty)}=\beta_{H}(y)_{[M,\infty)}$.
@@
To see this, let $C^{\prime\prime}$ be the component with the property that $\beta_{H}(L_{H^{\prime\prime}}(z^{\prime}))$ is backward asymptotic to $X_{C^{\prime\prime}}$. Then
@@
Indeed, $L_{H^{\prime\prime}}(\beta_{H}(L_{H^{\prime\prime}}(z^{\prime}))_{[M,\infty)}=L_{H^{\prime\prime}}(z^{\prime})_{[M,\infty)}=y_{[M,\infty)}$ by iii), which implies that
@@
by the choice of $M$, while the opposite inclusion follows from
@@
Since
@@
for all $j$ we have that $M(C^{\prime})\leq M(C^{\prime\prime})$. But it follows from (12) that $C^{\prime\prime}$ connects to $C$ and hence $M(C^{\prime\prime})\leq M(C^{\prime})$ by definition of $C^{\prime}$ and we find therefore that $M(C^{\prime\prime})=M(C^{\prime})$. It follows from this identity and (13) that $s_{H^{\prime\prime}}(z^{\prime}_{j})=s_{H^{\prime\prime}}(\beta_{H}(L_{H^{\prime\prime}}(z^{\prime}))_{j})$ for all $j$ ’close to -$\infty$’; a conclusion which implies the claim; (11). The inclusion $X_{C}\subseteq\overline{\beta_{H}(Y)}$ follows now because the path $\gamma$ is arbitrary and $\beta_{H}(Y)$ is $\sigma$-invariant. ∎
@@
We denote by $H^{\prime}$ the hereditary subgraph of $H^{\prime\prime}$ generated by the components $C$ of $H^{\prime\prime}$ with the property that there is a point $y\in Y$ such that $\beta_{H}(y)$ is backward asymptotic to $X_{C}$.
To simplify notation, set $L_{H^{\prime}}:=L_{H^{\prime\prime}}|_{H^{\prime}}$.
@@
commutes. Note that $L_{G}^{-1}(y)\in X_{G}^{L_{G}}\cap\kappa^{-1}(X_{G^{\prime}})$ for all $y\in Y$ since $\beta_{G}(Y)\subseteq X_{G^{\prime}}$ by definition. It follows from (15) that $\lambda^{\prime}(L_{G}^{-1}(y))\subseteq L_{H}^{-1}(\psi(y))$ and by symmetry that $\lambda^{\prime}(L_{G}^{-1}(y))=L_{H}^{-1}(\psi(y))$ for all $y\in Y$. Hence $\lambda(\beta_{G}(y))=\beta_{H}(\psi(y))$ for all $y\in Y$. In particular, $\lambda(\beta_{G}(y))=\phi^{\prime}(\beta_{G}(y))$ for all $y\in Y$ and hence, by Lemma 4.4, $\lambda|_{X_{C}}=\phi^{\prime}|_{X_{C}}$ for $C\in\mathcal{C}_{G}$. Consider then an $x\in X_{G^{\prime}}$ backward asymptotic to $X_{C},\ C\in\mathcal{C}_{G}$. Since $\lambda$ and $\phi^{\prime}$ agree on $X_{C}$ it follows that there is an $N\in\mathbb{Z}$ such that $\lambda(x)_{(-\infty,N]}=\phi^{\prime}(x)_{(-\infty,N]}$. Since $L_{H^{\prime}}$ is right-resolving and $L_{H^{\prime}}(\lambda(x))=\psi(L_{G^{\prime}}(x))=L_{H^{\prime}}(\phi^{\prime}(x))$, we find that $\lambda(x)=\phi^{\prime}(x)$. Since such $x$’s are dense in $X_{G^{\prime}}$ we conclude that $\lambda=\phi^{\prime}$. ∎

### v3

@@ 2.5. The terminal component in the future cover of an irreducible sofic shift
The existence of $\psi_{\mathbb{F}}$ follows from Corollary 2.24, Proposition 2.22 and Theorem 2.11. Let $\chi(Y)$ denote the set of elements $x\in X_{\mathbb{F}(Y)}$ for which $L_{\mathbb{F}(Y)}^{-1}(L_{\mathbb{F}(Y)}(x))=\{x\}$. It follows from Lemma 2.21 that $\mathbb{F}(Y)$ contains a finite path $\gamma$ such that every path in $\mathbb{F}(Y)$ with the label $L_{\mathbb{F}(Y)}(\gamma)$ terminates at the same vertex as $\gamma$. Since $\mathbb{F}(Y)$ is strongly connected every ray in $X_{\mathbb{F}(Y)}$ can be approximated by elements that contain $\gamma$ infinitely often to the left, and these elements are in $\chi(Y)$ because $(\mathbb{F}(Y),L_{\mathbb{F}(Y)})$ is right-resolving. Hence $\chi(Y)$ is dense $X_{\mathbb{F}(Y)}$.
If $\phi:X_{\mathbb{F}(Y)}\to X_{\mathbb{F}(Z)}$ is a conjugacy such that (10) commutes, $\phi(\chi(Y))=\chi(Z)$. If $\phi_{1}$ and $\phi_{2}$ are both conjugacies that make the diagram (10) commute, and $x\in\chi(Y)$, we have that $\phi_{i}(x)\in\chi(Z)$ and
$L_{\mathbb{F}(Z)}(\phi_{1}(x))=\psi(L_{\mathbb{F}(Y)}(x))=L_{\mathbb{F}(Z)}(\phi_{2}(x)),$
implying that $\phi_{1}(x)=\phi_{2}(x)$. Since $\chi(Y)$ is dense in $X_{\mathbb{F}(Y)}$ it follows that $\phi_{1}=\phi_{2}$. ∎
@@ 3. Additional properties of the future cover
Lemma 2.18 can now be improved to the following.
@@
Assume that $(G,L_{G})$ is a right-resolving and regular presentation of the sofic shift $Y$. It follows that there is a labeled graph homomorphism $\theta:(G,L_{G})\to(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ such that $\theta\left(\mathcal{R}(L_{G})\right)\subseteq\mathcal{R}(L_{\mathbb{K}(Y)})$, and $\theta(V_{G})$ is a hereditary subset of $V_{\mathbb{K}(Y)}$. If, furthermore, $(G,L_{G})$ is also follower-separated, $\theta$ is a labeled graph isomorphism of $(G,L_{G})$ onto the hereditary labeled subgraph $(\mathbb{K}(Y)_{\theta(V_{G})},L_{\mathbb{K}(Y)})$.
@@
By Lemma 2.18 and its proof there is a labeled-graph homomorphism $\theta:(G,L_{G})\to(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ such that $\theta(v)=F(L_{G}(x))$ where $x\in X_{G}$, $t_{G}(x_{(-\infty,-1]})=v$ and $f_{G}(v)=F(L_{G}(x))$. Let $F(\theta(v))\overset{a}{\to}F(z)$ be a labeled arrow in $\mathbb{K}(Y)$. Then $a$ is the first letter in an element $w\in F(L_{G}(x)))=f_{G}(v)$, and there is therefore a right-infinite path $\gamma$ in $G$ such that $s_{G}(\gamma)=v$ and $L_{G}(\gamma)=w$. The first arrow in $\gamma$ is an arrow $e\in E_{G}$ with $s_{G}(e)=v$ and $L_{G}(e)=a$. There is therefore an arrow $F(\theta(v))\overset{a}{\to}\theta(t_{G}(e))$ in $\mathbb{K}(Y)$. Since $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ is right-resolving this implies that $F(z)=\theta(t_{G}(e))$, proving that $\theta(V_{G})$ is a hereditary subset of $V_{\mathbb{K}(Y)}$. The remaining assertions follow now from Lemma 2.18.
@@
3.2. A universal property of the future cover $L_{\mathbb{K}(Y)}:X_{\mathbb{K}(Y)}\to Y$.
@@
Let $Y$ be an irreducible sofic shift. Let $X$ be an irreducible SFT and $\pi:X\to Y$ a right-closing factor map. There is a sliding block code $\rho:X\to X_{\mathbb{F}(Y)}$ such that
@@
Let $Y$ be an irreducible SFT and $\pi:X\to Y$ a factor map with $X$ an SFT. Then $\pi$ is isomorphic to the minimal right-resolving cover $L_{\mathbb{F}(Y)}:X_{\mathbb{F}(Y)}\to Y$ of $Y$ if and only if
@@
The following arguments are all implicitly contained in [BKM]. First, it is well-known from [BKM] and [LM] that the minimal right-resolving cover has all four properties, and hence so has any cover isomorphic to it. So assume that $\pi:X\to Y$ is a cover for which they all hold. We must show that $\pi$ is then isomorphic to the minimal right-resolving cover $L_{\mathbb{F}(Y)}:X_{\mathbb{F}(Y)}\to Y$ of $Y$. For this note that there are sliding block codes $\rho:X\to X_{\mathbb{F}(Y)}$ and $\rho^{\prime}:X_{\mathbb{F}(Y)}\to X$ such that
@@
both commute. An application of Lemma 9.1.13 in [LM] shows that $\rho$ and $\rho^{\prime}$ both maps doubly transitive points to doubly transitive points, so condition c) implies that $\rho\circ\rho^{\prime}=\operatorname{id}_{X_{\mathbb{F}(Y)}}$ and $\rho^{\prime}\circ\rho=\operatorname{id}_{X}$, first on doubly transitive points and then all over by continuity. ∎
@@
Let $(H,L_{H})$ be a regular right-resolving presentation of a sofic shift. For $x\in X_{H}$ and $N\in\mathbb{N}$ there is a $y\in\mathcal{R}(L_{H})$ such that $y_{i}=x_{i}$ for $i\geq-N$.
@@
Let $(H,L_{H})$ be a right-resolving presentation of a sofic shift. Then $(H,L_{H})$ is regular if and only if $\mathcal{R}(L_{H})$ is dense in $X_{H}$.
@@
Let $x\in\mathcal{R}(\pi)$ and consider an element $y\in\mathbb{U}(\pi^{\prime}(\rho(x)))=\mathbb{U}(\pi(x))$. There is then an element $z\in\mathbb{U}(x)$ such that $\pi(z)=y$. Since $\rho(z)\in\mathbb{U}(\rho(x))$ and $\pi^{\prime}(\rho(z))=y$, it follows that $\rho(x)\in\mathcal{R}(\pi^{\prime})$. ∎
@@ 4.1. Subset constructions
for $K\in X^{\#}$. In particular, we get from $(X,\phi)$ a new dynamical system $(X^{\#},\phi^{\#})$. More generally, when $(X,\phi)$ and $(Y,\phi^{\prime})$ are dynamical systems and $\pi:X\to Y$ is a map of dynamical systems in the sense that $\pi\circ\phi=\phi^{\prime}\circ\pi$, the set
@@
Consider two conjugacies $\psi^{\prime}:(X,\phi)\to(X^{\prime},\phi^{\prime})$ and $\psi:(Y,\mu)\to(Y^{\prime},\mu^{\prime})$ of dynamical systems, and factor maps $\pi:(X,\phi)\to(Y,\mu)$ and $\pi^{\prime}:(X^{\prime},\phi^{\prime})\to(Y^{\prime},\mu^{\prime})$ such that
@@
${X^{\pi}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi^{\prime\#}}$$\scriptstyle{\pi^{\#}}$${X^{\prime\pi^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi^{\prime\#}}$${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$${Y^{\prime}}$
@@ 4.2. Subset constructions for sofic shifts
To construct an inverse map $\kappa^{-1}:X_{H^{\prime\prime}}\to X_{H}^{L_{H}}$ to $\kappa$, let $x=(x_{i})_{i\in\mathbb{Z}}\in X_{H^{\prime\prime}}$. Since $(H,L_{H})$ is right-resolving, $s_{H^{\prime\prime}}(x_{i})\geq\#s_{H^{\prime\prime}}(x_{i+1})$, and we choose $N\in\mathbb{N}$ such that $s_{H^{\prime\prime}}(x_{i})=\#s_{H^{\prime\prime}}(x_{N})$ for all $i\leq N$. For each $v\in s_{H^{\prime\prime}}(x_{N})$ there is a unique ray $\gamma^{v}\in X_{H}$ such that $s_{H}(\gamma^{v}_{N})=v$, $s_{H}(\gamma^{v}_{i})\in s_{H^{\prime\prime}}(x_{i})$ for all $i\in\mathbb{Z}$ and $L_{H}(\gamma^{v})=L_{H^{\prime\prime}}(x)$. Setting
@@
Thus $({G^{\prime\prime}},L_{G^{\prime\prime}})$ is a weakly canonical cover when $(G,L_{G})$ is. However, the conjugacy $\phi^{\prime\prime}$ may not be unique and there is no analogue to the right-inverse for the labeling map as in Krieger’s theorem. We continue therefore to identify a labeled subgraph of $(G^{\prime\prime},L_{G^{\prime\prime}})$ which does give rise to a strongly canonical cover.
@@
(11) $L_{H^{\prime\prime}}\circ\beta_{H}=\operatorname{id}_{Y}.$
Like the map $\alpha_{Y}$ considered in Section 2.2 the map $\beta_{H}$ commutes with the shift but is not continuous.
@@
Let $y\in Y$ and let $C$ be the component in $H^{\prime\prime}$ for which there is a $N\in\mathbb{Z}$ such that $\beta_{H}(y)_{j}\in C$ for all $j\leq N$. It follows that there is a $R<N$ such that if $z\in X_{H^{\prime\prime}}$ and $L_{H^{\prime\prime}}(z)_{[R,\infty)}=y_{[R,\infty)}$, then $s_{H^{\prime\prime}}(z_{N})\subseteq s_{H^{\prime\prime}}(\beta_{H}(y)_{N})$. (If not, a compactness argument would produce an element $z^{\prime}\in X_{H^{\prime\prime}}$ such that $L_{H^{\prime\prime}}(z^{\prime})=y$ and $s_{H^{\prime\prime}}(z^{\prime}_{N})>\#s_{H^{\prime\prime}}(\beta_{H}(y)_{N})$. A contradiction.) Let $C^{\prime}$ be a component which connects to $C$ and has the largest multiplicity among the components that connect to $C$. Let $\gamma$ be a finite path in $H^{\prime\prime}_{C}$. We can then choose an element $z^{\prime}\in X_{H^{\prime\prime}}$ with the following properties:
@@
    there is a $R^{\prime}<R$ such that $z^{\prime}_{[R^{\prime},R]}$ is a path in $H^{\prime\prime}_{C}$ with $\gamma\subseteq z^{\prime}_{[R^{\prime},R]}$, and
@@
    $z^{\prime}_{[R,\infty)}=\beta_{H}(y)_{[R,\infty)}$.
@@
To see this, we use that
@@
Indeed, $L_{H^{\prime\prime}}(\beta_{H}(L_{H^{\prime\prime}}(z^{\prime})))_{[R,\infty)}=L_{H^{\prime\prime}}(z^{\prime})_{[R,\infty)}=y_{[R,\infty)}$ by iii), which implies that
@@
by the choice of $R$, while the opposite inclusion follows from
@@
Let $C^{\prime\prime}$ be the component in $H^{\prime\prime}$ with the property that $\beta_{H}(L_{H^{\prime\prime}}(z^{\prime}))$ is backward asymptotic to $X_{C^{\prime\prime}}$. Since
@@
for all $j$ and $z^{\prime}$ is backward asymptotic to $X_{C^{\prime}}$ while $\beta_{H}(L_{H^{\prime\prime}}(z^{\prime}))$ is backward asymptotic to $X_{C^{\prime\prime}}$, we have that $M(C^{\prime})\leq M(C^{\prime\prime})$. But it follows from (13) that $C^{\prime\prime}$ connects to $C$ and hence $M(C^{\prime\prime})\leq M(C^{\prime})$ by definition of $C^{\prime}$, and we find therefore that $M(C^{\prime\prime})=M(C^{\prime})$. It follows from this identity and (14) that $s_{H^{\prime\prime}}(z^{\prime}_{j})=s_{H^{\prime\prime}}(\beta_{H}(L_{H^{\prime\prime}}(z^{\prime}))_{j})$ for all $j$ ’close to -$\infty$’; a conclusion which implies the claim; (12). The inclusion $X_{C}\subseteq\overline{\beta_{H}(Y)}$ follows now from (12) because the path $\gamma$ is arbitrary and $\beta_{H}(Y)$ is $\sigma$-invariant. ∎
@@
We denote by $H^{\prime}$ the hereditary subgraph of $H^{\prime\prime}$ generated by the components $C$ of $H^{\prime\prime}$ for which there is a point $y\in Y$ such that $\beta_{H}(y)$ is backward asymptotic to $X_{C}$.
To simplify notation, set $L_{H^{\prime}}:=L_{H^{\prime\prime}}|_{H^{\prime}}$. Since $\beta_{H}(Y)\subseteq X_{G^{\prime}}$ by definition of $G^{\prime}$, it follows from (11) that $L_{X_{G^{\prime}}}(X_{G^{\prime}})=Y$, implying that $(G^{\prime},L_{G^{\prime}})$ is a presentation of $Y$.
@@
commutes. Note that $L_{G}^{-1}(y)\in X_{G}^{L_{G}}\cap\kappa^{-1}(X_{G^{\prime}})$ for all $y\in Y$ since $\beta_{G}(Y)\subseteq X_{G^{\prime}}$. It follows from (16) that $\lambda^{\prime}(L_{G}^{-1}(y))\subseteq L_{H}^{-1}(\psi(y))$ and by symmetry that $\lambda^{\prime}(L_{G}^{-1}(y))=L_{H}^{-1}(\psi(y))$ for all $y\in Y$. Hence $\lambda(\beta_{G}(y))=\beta_{H}(\psi(y))$ for all $y\in Y$. In particular, $\lambda(\beta_{G}(y))=\phi^{\prime}(\beta_{G}(y))$ for all $y\in Y$ and hence, by Lemma 4.4, $\lambda|_{X_{C}}=\phi^{\prime}|_{X_{C}}$ for $C\in\mathcal{C}_{G}$. Consider then an $x\in X_{G^{\prime}}$ backward asymptotic to $X_{C},\ C\in\mathcal{C}_{G}$. Since $\lambda$ and $\phi^{\prime}$ agree on $X_{C}$ it follows that there is an $N\in\mathbb{Z}$ such that $\lambda(x)_{(-\infty,N]}=\phi^{\prime}(x)_{(-\infty,N]}$. Since $L_{H^{\prime}}$ is right-resolving and $L_{H^{\prime}}(\lambda(x))=\psi(L_{G^{\prime}}(x))=L_{H^{\prime}}(\phi^{\prime}(x))$, we find that $\lambda(x)=\phi^{\prime}(x)$. Since such $x$’s are dense in $X_{G^{\prime}}$ we conclude that $\lambda=\phi^{\prime}$. ∎